OpenStudy (anonymous):

Please help! Need to find eccentricity of this Hyperbola:

OpenStudy (anonymous):

\[5x ^{2}-9y ^{2}=45\]

OpenStudy (anonymous):

What are a and b ?

OpenStudy (anonymous):

oh gosh wait i can divide both sides by 45 to get it to equal 1 and find a and b right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so a is 3 and b is square root 5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and i can use c^2=a^2+b^2 formula to find c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so e=7/3?

OpenStudy (anonymous):

recheck your answer

OpenStudy (anonymous):

oh oops its square root 14/3?

OpenStudy (anonymous):

That's right !

OpenStudy (anonymous):

ok thanks!